# Optimal and Practical Batched Linear Bandit Algorithm

### v1

@@ 7.1 Experimental Setup
In this section, we empirically evaluate the performance of the BLAE algorithm. We measure the cumulative regret over $T=100,000$ time steps. The $K$ arms and $\theta^{*}$ are randomly sampled from a $d$-dimensional uniform distribution. The experiment is repeated 10 times, with the arms being independently resampled from the uniform distribution in each run. We conduct experiments with the following $(K,d)$ parameter pairs $(K,d)=(50,5),(50,10),$ $(100,5),(100,10),(200,5),(200,10),(400,5),(400,10)$.
@@
Algorithms proposed by Ruan et al. (2021); Hanna et al. (2023); Zhang et al. (2025) incur significant computational overhead; their runtimes are reported in Section E, as explanation for their empirical inefficiency is presented in Section 6. Accordingly, we omit regret plots for these methods.
@@ 7.2 Experimental Results
Among recently proposed methods, $\texttt{E}^{4}$ (Ren et al., 2024) is shown to attain minimal batch complexity (Figure 1). This advantage, however, comes at a cost in regret performance: the algorithm records the highest cumulative regret and shows marked instability. Minimizing the number of batches is therefore of limited value if the resulting regret is unsustainable. Across 80 independent experiments, $\texttt{E}^{4}$ shows substantial performance variability due to its aggressive arm-elimination policy and sensitivity to batch size. As noted in Section 6, its elimination strategy fails to respect independence in the batched setting, occasionally discarding the optimal arm prematurely. In addition, the algorithm’s sensitivity to the hyperparameter $\gamma$ often causes the third batch to consume almost the entire horizon, leaving little opportunity for exploitation. Our evaluation used the official implementation released by Ren et al. (2024).
@@ Appendix A Proof of Lemma 1
The proof of Lemma 1 relies on the following two technical lemmas. Here, we assume that subgaussian random variables have zero mean.
@@
The matrix $H_{\ell}$ is independent of the noise terms $\{\eta_{s}\}_{s=t_{\ell-1}+1}^{t_{\ell}}$. Furthermore, these $1$-subgaussian random variables $\{\eta_{s}\}_{s=t_{\ell-1}+1}^{t_{\ell}}$ are mutually independent. According to Lemma 5, the term
@@
Applying the concentration inequality from Lemma 4 to this subgaussian random variable yields
@@
$\sum_{s=t_{\ell-1}+1}^{t_{\ell}}\langle x,H_{\ell}^{-1}x_{s}\rangle^{2}\leq x^{T}H_{\ell}^{-1}(\sum_{s=t_{\ell-1}+1}^{t_{\ell}}x_{s}x_{s}^{T}+\lambda I)H_{\ell}^{-1}x=x^{T}H_{\ell}^{-1}H_{\ell}H_{\ell}^{-1}x=\|x\|_{H_{\ell}^{-1}}^{2}\;.$
This completes the proof of Lemma 1:
@@ Appendix B Proof of Lemma 2
The proof of Lemma 2 relies on the following five technical lemmas.
@@
Let $\zeta\in(0,1).$ There exists a set $C_{\zeta}\subset S^{d-1}$ with cardinality $|C_{\zeta}|\leq\sqrt{2\pi d}\,(\zeta^{2}-\frac{\zeta^{4}}{4})^{-\frac{d-1}{2}}$ such that for any $x\in S^{d-1}$, there exists a $y\in C_{\zeta}$ satisfying $\|x-y\|\leq\zeta$, where $S^{d-1}\coloneqq\{x\in\mathbb{R}^{d}:\|x\|=1\}$ and $d\geq 2$.
@@
$N(\zeta)\coloneqq\min\{|C|:C\text{ is an $\zeta$-cover of }S^{d-1}\},\quad M(\zeta)\coloneqq\max\{|P|:P\text{ is an $\zeta$-packing of }S^{d-1}\}\;.$
@@
From (4) and Lemma 6, and the arbitrariness of $\zeta$-packing $P$, we obtain
@@
Using the monotonicity of $t\rightarrow(1-t)^{-\frac{1}{2}}$ on $(0,\frac{3}{4})$, along with $0<\zeta^{2}(1-\frac{\zeta^{2}}{4})<\frac{3}{4}$ for $\zeta\in(0,1)$, and applying Lemma 7, we derive
@@
[…]
@@
After the $\ell$-th batch $(\ell\geq 1)$, the following bound holds for any $10<\zeta<$ 1 and $d\geq 2$ with probability at least $1-\delta$
$\|\hat{\theta}_{{\ell}}-\theta^{*}\|_{H_{{\ell}}}\leq\frac{\sqrt{(d-1)\log(\frac{4}{4\zeta^{2}-\zeta^{4}})+2\log(\frac{\sqrt{2\pi d}}{\delta})}+\sqrt{\lambda}}{1-\zeta}\;.$
@@
Let $C_{\zeta}\subset S^{d-1}$ be the $\zeta$-covering set given in Lemma 8, and define the event
@@
Using the property $\|H_{\ell}^{\frac{1}{2}}x\|_{H_{\ell}^{-1}}=\|x\|=1$ for all $x\in S^{d-1}$, and by applying Lemma 1, we have $\mathbb{P}(E)\leq\delta$.
@@
Therefore, we prove Lemma 9:
@@
From this point onward, for convenience, we define $M_{\ell}$ as follows
@@
Here, $\zeta\in(0,1)$ is a tunable parameter. While it is possible to optimize over $\zeta$, we fix $\zeta=0.5$ for simplicity, as this choice influences the results only up to a constant factor. Nevertheless, we retain $\zeta$ as a symbolic variable in the following analysis to maintain generality. Readers may safely substitute $\zeta=0.5$ throughout.
@@
Using the Cauchy-Schwarz inequality and Lemma 9, the following result holds for all $x,y\in\mathcal{A}_{\ell-1}$ with probability at least $1-\frac{\delta}{2}$
@@
Next, by applying Lemma 1 to the $\binom{|\mathcal{A}_{\ell-1}|}{2}$ pairs of $x,y\in\mathcal{A}_{\ell-1}$ such that $\langle x-y,\hat{\theta}_{\ell}-\theta^{*}\rangle\geq 0$, we know that for all $x,y\in\mathcal{A}_{\ell-1}$, the following holds with probability at least $1-\frac{\delta}{2}$
@@
Combining the result from (5) and (6), we prove Lemma 10:
@@
Let $E_{\ell}\coloneqq\{|\langle x-y,\hat{\theta}_{\ell}-\theta^{*}\rangle|\leq\varepsilon_{\ell}(\delta),\;\forall x,y\in\mathcal{A}_{\ell-1}\}$. By Lemma 10, we have
@@
[…]
Therefore, we prove Lemma 2:
@@ Appendix C Proof of Lemma 3
The proof of Lemma 3 relies on the following four technical lemmas.
@@
Therefore, we prove Lemma 11:
@@
Using Lemma 11, for any symmetric matrix $U\in\mathbb{R}^{d\times d}$, the Hessian of $g(X)$ is computed as follows
@@
$g(V_{\ell}(\lambda\pi_{\ell}+(1-\lambda)\widetilde{\pi}_{\ell}))=g(\lambda V_{\ell}(\pi_{\ell})+(1-\lambda)V_{\ell}(\widetilde{\pi}_{\ell}))\geq\lambda g(V_{\ell}(\pi_{\ell}))+(1-\lambda)g(V_{\ell}(\widetilde{\pi}_{\ell}))\;.$
@@
Therefore, we prove Lemma 13:
@@
Using Lemma 11, we obtain
@@
Since $\mathcal{D}_{\ell}$ is compact and $f(\pi_{\ell})$ is concave by Lemma 12, Lemma 13 implies
@@
00 0 $\geq\|x^{(i)}\|_{V_{\ell}(\pi_{\ell}^{\prime})^{-1}}^{2}-\sum_{j=1}^{|\mathcal{A}_{\ell-1}|}\|x^{(j)}\|_{V_{\ell}(\pi_{\ell}^{\prime})^{-1}}^{2}\pi_{\ell j}^{\prime}$
@@
Therefore, we prove Lemma 3:
@@ Appendix D Proof of Theorem 1
By Lemma 2, this nice event occurs with high probability
@@ E.1 Runtime Comparison Analysis
We provide a runtime comparison with batched linear bandit algorithms in Ruan et al. (2021); Hanna et al. (2023); Zhang et al. (2025). The experiments were conducted with $K=50$ arms and a feature dimension of $d=5$, where all arms and the unknown parameter $\theta^{*}$ were sampled from a normal distribution. Under the same experimental setting, the BLAE algorithm requires only $7.3$ seconds for 10 runs when $T=$ 100,000. In contrast, as shown in Table 2, the algorithm in Ruan et al. (2021) requires $4642.2$ seconds for just $T=5000$, while the algorithm in Zhang et al. (2025) takes $7178.1$ seconds for $T=5000$, demonstrating their limitation in large-scale problems. Furthermore, the algorithm in Hanna et al. (2023) faces severe computational challenges due to the construction of $\Theta^{\prime}$ as mentioned in Section 6. The size of $\Theta^{\prime}$ is $\mathcal{\mathcal{O}}(T^{d})$, meaning that for $T=1000$, $\Theta^{\prime}$ would contain approximately $10^{15}$ vectors, making the algorithm infeasible even for moderate values of $T$. Due to these computational constraints, we were unable to measure its runtime. From this experiment, we conclude that the batched linear bandit algorithms proposed in Ruan et al. (2021); Zhang et al. (2025); Hanna et al. (2023) are impractical for our setting.
Table 2: Runtime (seconds) results for the normal distribution setting with $K=50$ and $d=5$.

### v2

@@ 7.1 Experimental Setup
In this section, we empirically evaluate the performance of the BLAE algorithm. We measure the cumulative regret over $T=100,000$ time steps. The $K$ arms and $\theta^{*}$ are randomly sampled from a $d$-dimensional uniform distribution. The experiment is repeated 10 times, with the arms being independently resampled from the uniform distribution in each run. We conduct experiments with the following $(K,d)$ parameter pairs $(K,d)=(50,5),(50,10),$ (100,5),(100,10),(200,5),(200,10),(400,5),(400,10)(100,5),(100,10),(200,5),(200,10),(400,5),(400,10).
@@
Algorithms proposed by Ruan et al. (2021); Hanna et al. (2023); Zhang et al. (2025) incur significant computational overhead; their runtimes are reported in Section E, as explanation for their empirical inefficiency is presented in Section˜6. Accordingly, we omit regret plots for these methods.
@@ 7.2 Experimental Results
Among recently proposed methods, $\texttt{E}^{4}$ (Ren et al., 2024) is shown to attain minimal batch complexity (Figure 1). This advantage, however, comes at a cost in regret performance: the algorithm records the highest cumulative regret and shows marked instability. Minimizing the number of batches is therefore of limited value if the resulting regret is unsustainable. Across 80 independent experiments, $\texttt{E}^{4}$ shows substantial performance variability due to its aggressive arm-elimination policy and sensitivity to batch size. As noted in Section˜6, its elimination strategy fails to respect independence in the batched setting, occasionally discarding the optimal arm prematurely. In addition, the algorithm’s sensitivity to the hyperparameter $\gamma$ often causes the third batch to consume almost the entire horizon, leaving little opportunity for exploitation. Our evaluation used the official implementation released by Ren et al. (2024).
@@ Appendix A Proof of Lemma 1
The proof of Lemma˜1 relies on the following two technical lemmas. Here, we assume that subgaussian random variables have zero mean.
@@
The matrix $H_{\ell}$ is independent of the noise terms $\{\eta_{s}\}_{s=t_{\ell-1}+1}^{t_{\ell}}$. Furthermore, these $1$-subgaussian random variables $\{\eta_{s}\}_{s=t_{\ell-1}+1}^{t_{\ell}}$ are mutually independent. According to Lemma˜5, the term
@@
Applying the concentration inequality from Lemma˜4 to this subgaussian random variable yields
@@
$\sum_{s=t_{\ell-1}+1}^{t_{\ell}}\langle x,H_{\ell}^{-1}x_{s}\rangle^{2}\leq x^{T}H_{\ell}^{-1}\left(\sum_{s=t_{\ell-1}+1}^{t_{\ell}}x_{s}x_{s}^{T}+\lambda I\right)H_{\ell}^{-1}x=x^{T}H_{\ell}^{-1}H_{\ell}H_{\ell}^{-1}x=\|x\|_{H_{\ell}^{-1}}^{2}\;.$
This completes the proof of Lemma˜1:
@@ Appendix B Proof of Lemma 2
The proof of Lemma˜2 relies on the following five technical lemmas.
@@
For any $\zeta\in(0,1)$, there exists a finite cover $C_{\zeta}\subset S^{d-1}$ such that for all $x\in S^{d-1}$, there exists $y\in C_{\zeta}$ with $\|x-y\|\leq\zeta$, and $|C_{\zeta}|\leq\sqrt{2\pi d}\cdot(\zeta^{2}-\zeta^{4}/4)^{-\frac{d-1}{2}}$. Here, $S^{d-1}\coloneqq\{x\in\mathbb{R}^{d}:\|x\|=1\}$ denotes the (d1)(d-1)-dimensional unit sphere in $\mathbb{R}^{d}$, and $d\geq 2$.
@@
[…]
@@
From (4) and Lemma˜6, and the arbitrariness of $\zeta$-packing $P$, we obtain
@@
Using the monotonicity of $t\rightarrow(1-t)^{-\frac{1}{2}}$ on $(0,\frac{3}{4})$, along with $0<\zeta^{2}(1-\frac{\zeta^{2}}{4})<\frac{3}{4}$ for $\zeta\in(0,1)$, and applying Lemma˜7, we derive
@@
$=\frac{2B(\frac{d-1}{2},\frac{1}{2})}{\int_{0}^{\zeta^{2}(1-\frac{\zeta^{2}}{4})}t^{\frac{d-3}{2}}(1-t)^{-\frac{1}{2}}dt}$
$\leq\frac{2B(\frac{d-1}{2},\frac{1}{2})}{\int_{0}^{\zeta^{2}(1-\frac{\zeta^{2}}{4})}t^{\frac{d-3}{2}}dt}$
@@
After the $\ell$-th batch $(\ell\geq 1)$, the following bound holds for any $0<\zeta<1$ and $d\geq 2$ with probability at least $1-\delta$
$\|\hat{\theta}_{{\ell}}-\theta^{*}\|_{V_{{\ell}}}\leq\frac{\sqrt{(d-1)\log(\frac{4}{4\zeta^{2}-\zeta^{4}})+2\log(\frac{\sqrt{2\pi d}}{\delta})}+\sqrt{\lambda}}{1-\zeta}\;.$
@@
Let $C_{\zeta}\subset S^{d-1}$ be the $\zeta$-covering set given in Lemma˜8, and define the event
@@
Using the property $\|H_{\ell}^{\frac{1}{2}}x\|_{H_{\ell}^{-1}}=\|x\|=1$ for all $x\in S^{d-1}$, and by applying Lemma˜1, we have $\mathbb{P}(E)\leq\delta$.
@@
Therefore, we prove Lemma˜9:
@@
Here, $\zeta\in(0,1)$ is a tunable parameter. Although one could optimize over $\zeta$, we fix $\zeta=0.5$ for simplicity, as this choice affects the results only up to a constant factor. We nevertheless retain $\zeta$ as a symbolic variable in the subsequent analysis to preserve generality; readers may safely substitute $\zeta=0.5$ throughout. For convenience, we define
@@
Using the Cauchy-Schwarz inequality and Lemma˜9, the following result holds for all $x,y\in\mathcal{A}_{\ell-1}$ with probability at least $1-\frac{\delta}{2}$
@@
Next, by applying Lemma˜1 to the $\binom{|\mathcal{A}_{\ell-1}|}{2}$ pairs of $x,y\in\mathcal{A}_{\ell-1}$ such that $\langle x-y,\hat{\theta}_{\ell}-\theta^{*}\rangle\geq 0$, we know that for all $x,y\in\mathcal{A}_{\ell-1}$, the following holds with probability at least $1-\frac{\delta}{2}$
@@
Combining the result from (5) and (6), we prove Lemma˜10:
@@
Let $E_{\ell}\coloneqq\{|\langle x-y,\hat{\theta}_{\ell}-\theta^{*}\rangle|\leq\varepsilon_{\ell}(\delta),\;\forall x,y\in\mathcal{A}_{\ell-1}\}$. By Lemma˜10, we have
@@
$\mathbb{P}(E^{c})=\mathbb{P}(\overset{B}{\underset{\ell=1}{\cup}}E_{\ell}^{c})\leq\sum_{\ell=1}^{B}\mathbb{P}(E_{\ell}^{c})\leq\frac{B\delta}{1+\lceil\log_{2}\log_{2}T\rceil}\leq\delta\;.$
Therefore, we prove Lemma˜2:
@@ Appendix C Proof of Lemma 3
The proof of Lemma˜3 relies on the following four technical lemmas.
@@
Therefore, we prove Lemma˜11:
@@
Using Lemma˜11, for any symmetric matrix $U\in\mathbb{R}^{d\times d}$, the Hessian of $g(X)$ is computed as follows
@@
$).g(V_{\ell}(\lambda\pi_{\ell}+(1-\lambda)\widetilde{\pi}_{\ell}))=g(\lambda V_{\ell}(\pi_{\ell})+(1-\lambda)V_{\ell}(\widetilde{\pi}_{\ell}))\geq\lambda g(V_{\ell}(\pi_{\ell}))+(1-\lambda)g(V_{\ell}(\widetilde{\pi}_{\ell}$))\;.
@@
Therefore, we prove Lemma˜13:
@@
Using Lemma˜11, we obtain
@@
Since $\mathcal{D}_{\ell}$ is compact and $f(\pi_{\ell})$ is concave by Lemma˜12, Lemma˜13 implies
@@
$0$ $\geq\|x^{(i)}\|_{V_{\ell}(\pi_{\ell}^{\prime})^{-1}}^{2}-\sum_{j=1}^{|\mathcal{A}_{\ell-1}|}\|x^{(j)}\|_{V_{\ell}(\pi_{\ell}^{\prime})^{-1}}^{2}\pi_{\ell j}^{\prime}$
@@
Therefore, we prove Lemma˜3:
@@ Appendix D Proof of Theorem 1
By Lemma˜2, this nice event occurs with high probability
@@ E.1 Runtime Comparison Analysis
We provide a runtime comparison with batched linear bandit algorithms in Ruan et al. (2021); Hanna et al. (2023); Zhang et al. (2025). The experiments were conducted with $K=50$ arms and a feature dimension of $d=5$, where all arms and the unknown parameter $\theta^{*}$ were sampled from a normal distribution. Under the same experimental setting, the BLAE algorithm requires an average of $0.73$ seconds over 10 runs when $T=$ 100,000. In contrast, as shown in Table 2, the algorithm in Ruan et al. (2021) requires an average of $464.2$ seconds for $T=5000$, while the algorithm in Zhang et al. (2025) takes an average of $717.8$ seconds for $T=5000$, demonstrating their limitation in large-scale problems. Furthermore, the algorithm in Hanna et al. (2023) faces severe computational challenges due to the construction of $\Theta^{\prime}$ as mentioned in Section 6. The size of $\Theta^{\prime}$ is $\mathcal{\mathcal{O}}(T^{d})$, meaning that for $T=1000$, $\Theta^{\prime}$ would contain approximately $10^{15}$ vectors, making the algorithm infeasible even for moderate values of $T$. Due to these computational constraints, we were unable to measure its runtime. From this experiment, we conclude that the batched linear bandit algorithms proposed in Ruan et al. (2021); Zhang et al. (2025); Hanna et al. (2023) are impractical for our setting.
Table 2: Average runtime (seconds) over 10 runs for the normal distribution setting with $K=50$ and $d=5$.